Help!!!

Find the radius of a circle with an area of 314 sq mm.

Thank you

Answer by checkley75(3666)

AREA=PI*R^2
314=3.14*R^2
3.14R^2=314
R^2=314/3.14
R^2=100 TAKE THE SQUARE OF BOTH SIDES 7 WE GET
R=10 MM